OpenStudy (anonymous):

Write an inequality that represents the statement "x is greater than –3 and less than or equal to 4." @Dobby1

OpenStudy (dobby1):

is there a reason you mentioned me just curious

OpenStudy (anonymous):

-3 < x ≤ 4

OpenStudy (anonymous):

no i just need help with this question on my homework

OpenStudy (dobby1):

ok this question is easy I am going to give an example that I will make up

OpenStudy (dobby1):

and @Shirobara plz do not just answer question you must give explanations atleast

OpenStudy (anonymous):

okay thanks

OpenStudy (dobby1):

ok so the question I am going to make up is x is greater that 50 but less that 100

OpenStudy (anonymous):

see i dont under stand

OpenStudy (dobby1):

it is not to hard to solve when they say greater that they want this sign > when they want less than they want this sign < \[when they want less than or equal \to they want this sign \] \[when they want greater than or equal \to they want this sign \]

OpenStudy (anonymous):

<

OpenStudy (dobby1):

ok apparently you can not type in the equation box

OpenStudy (dobby1):

I will show them in a list

OpenStudy (dobby1):



OpenStudy (dobby1):

top left corner is a lessthan sign bottom left corner is a greaterthan sign to right corner is a lessthan or equal too sign bottom right corner is a greater than or equal to sign

OpenStudy (dobby1):

are you following

OpenStudy (dobby1):

@jeremiah94278

OpenStudy (dobby1):

Hey @jeremiah94278 you still there?

OpenStudy (e.mccormick):

\(\text{when they want greater than or equal} \to \text{they want this sign}\) `\(\text{when they want greater than or equal} \to \text{they want this sign}\)`

OpenStudy (dobby1):

see it did not work for you either i was trying to put the sign in there

OpenStudy (anonymous):

oh okay

OpenStudy (dobby1):

are youfollowing @jeremiah94278

OpenStudy (e.mccormick):

Want the signs? \(\ge \le \ne\) ? I don't use the editor. I just type them. Ah, he is back!

OpenStudy (dobby1):

yeah ikr lol

OpenStudy (dobby1):

ok well back to the question I posted for you since it says greater than 50 but and less than 100 we would put 50<x<100 because it is greater than 50 but less than 100

OpenStudy (dobby1):

Are you following @jeremiah94278

OpenStudy (e.mccormick):

Here are the notes Dobby was doing: \( \le\) less than and equal to \(<\) less than \(\ge\) greater than and equal to \(>\) greater than \(\ne\) not equal to \(=\) equal \(\le\), \(\ge\), and \(\ne\) are dashed lines or circles (empty center) on graphs. \(<\), \(>\), and \(=\) are solid lines or dots on graphs.